# Attention-Guided Contrastive Role Representations for Multi-Agent Reinforcement Learning

### 08 Dec 2023 v1

@@ Abstract
Real-world multi-agent tasks usually involve dynamic team composition with the emergence of roles, which should also be a key to efficient cooperation in multi-agent reinforcement learning (MARL). Drawing inspiration from the correlation between roles and agent’s behavior patterns, we propose a novel framework of Attention-guided COntrastive Role representation learning for MARL (ACORM) to promote behavior heterogeneity, knowledge transfer, and skillful coordination across agents. First, we introduce mutual information maximization to formalize role representation learning, derive a contrastive learning objective, and concisely approximate the distribution of negative pairs. Second, we leverage an attention mechanism to prompt the global state to attend to learned role representations in value decomposition, implicitly guiding agent coordination in a skillful role space to yield more expressive credit assignment. Experiments and visualizations on challenging StarCraft II micromanagement tasks demonstrate the state-of-the-art performance of our method and its advantages over existing approaches. Our code is available at https://github.com/NJU-RL/ACORM.
Keywords Multi-agent reinforcement learning  $\cdot$ Role representation  $\cdot$ Contrastive learning  $\cdot$ Attention mechanism
@@ 1 Introduction
Cooperative multi-agent reinforcement learning (MARL) aims to coordinate a system of agents towards optimizing accumulated global rewards in a dynamically shifting environment [1], and has witnessed significant prospects in various applications, such as autonomous vehicles [2], traffic signal control [3], smart grid [4], robotics [5], and social science [6]. Training reliable control policies for coordinating such systems remains a major challenge. The framework of centralized training with decentralized execution (CTDE) [7] hybrids the merits of independent Q-learning [8, 9] and joint action learning [10, 11], and becomes a compelling paradigm that exploits the centralized training opportunity for training fully decentralized policies [12]. Subsequently, numerous popular algorithms are proposed, including VDN [13], QMIX [14], MAAC [15], and MAPPO [16].
Sharing policy parameters is crucial in scaling these algorithms to a large number of agents with accelerated cooperation learning [17]. However, it is widely observed that agents tend to acquire homogeneous behaviors, which might hinder diversified exploration and sophisticated coordination [18]. Some methods [19, 20, 21] attempt to promote individualized behaviors by distinguishing each agent from the others, while they often neglect the prospect of effective team composition with implicit task allocation. Real-world multi-agent tasks usually involve dynamic team composition with the emergence of roles [22, 23]. [^1] Taking the football game [24] as an example, the midfielders are primarily responsible for delivering the ball to the forwards to coordinate shots on goal in the offensive phase, while they need to drop back and join the defenders to block passing lanes on the defensive. Early works introduce the role concept into multi-agent systems [25, 26, 27], while they usually require prior domain knowledge to pre-define role responsibilities. Recently, ROMA [28] learns emergent roles conditioned solely on current observations, and RODE [29] associates each role with a fixed subset of the joint action space. COPA [30] allows dynamic role allocation via distributing a global view of team composition to each agent during execution. Some works decompose the task into a set of skills [31], subtasks [32, 33], or groups [34] with a hierarchical structure for control. Overall, existing role-based methods still suffer from several deficiencies, such as insufficient characterization of complex behaviors for role emergence, neglect of evolving team dynamics, or relaxation of the CTDE constraint.
@@
ACORM is fully compatible with CTDE methods, and we realize ACORM on top of two popular MARL algorithms, QMIX [14] and MAPPO [16], benchmarked on challenging StarCraft multi-agent challenge (SMAC) [35] and Google research football (GRF) [24] environments. Experiments demonstrate that ACORM achieves state-of-the-art performance on most scenarios. Visualizations of learned role representations, heterogeneous behavior patterns, and attentional value decomposition shed further light on our advantages. Ablation studies confirm that ACORM promotes higher coordination capacity by virtue of contrastive role representation learning and attention-guided credit assignment, respectively, even if agents have the same innate characteristics. In summary, our contributions are threefold:
@@ 2 Method
In this section, we present the ACORM framework. We consider cooperative multi-agent tasks formulated as a Dec-POMDP [36], $G=\langle I,S,A,P,R,\Omega,O,n,\gamma\rangle$, where $I$ is a finite set of $n$ agents, $s\in S$ is the global state, and $\gamma\in[0,1)$ is the discount factor. At each time step, each agent $i$ draws an observation $o_{i}\in O$ from $\Omega(s,i)$ and selects a local action $a_{i}\in A$. After executing the joint action $\bm{a}=[a_{1},...,a_{n}]^{\top}\in A^{n}$, the system transitions to a next state $s^{\prime}$ according to $P(s^{\prime}|s,\bm{a})$ and receives a reward $r=R(s,\bm{a})$ shared by all agents.
@@
Figure 1 illustrates ACORM based on QMIX, and Appendix B presents the corresponding algorithm pseudocode. Moreover, the extension to MAPPO is given in Appendix C. ACORM consists of individual Q-networks in Fig. 1(b) and a mixing network in Fig. 1(a). We introduce mutual information maximization to formalize role representation learning, and derive a contrastive learning objective that optimizes agent embeddings $\{e^{t}_{i}\}_{i=1}^{n}$ in a self-supervised way to acquire contrastive role representations $\{z_{i}^{t}\}_{i=1}^{n}$, which is shown in Fig. 1(c) and will be introduced in detail in Section 2.1. In value decomposition, we employ multi-head attention (MHA) to prompt the global state to attend to learned role patterns, guiding skillful agent coordination in the high-level role space for facilitating expressive credit assignment, as described in Fig. 1(d) and in Section 2.2.
@@ 2.1 Contrastive Role Representations
In practice, directly optimizing mutual information is intractable. Inspired by the noise contrastive estimation (InfoNCE) [37] in the literature of contrastive learning [38, 39], we derive a lower bound of Eq. (1) with the following theorem.
@@
Negative Pairs Generation. Periodically, we partition all $n$ agents into $K$ clusters $\{C_{j}\}_{j=1}^{K}$ according to agent embeddings. [^4] In this paper, we simply use K-means [40] based on Euclidean distances between agent embeddings. Moreover, it can be easily extended to more complex clustering methods such as Gaussian mixture models [41]. In Appendix D, we conduct the hyperparameter analysis about the influence of different $K$ values and how to determine the number of clusters automatically. Naturally, we encourage role representations from the same cluster to stay close to each other, while differing from agents in other clusters. For agent $i$, we denote its role representation $z_{i}$ as the query $q$, and role representations of other agents as the keys $\mathbb{K}=\{z_{1},...,z_{n}\}\backslash z_{i}$. Points from the same cluster as the query, $i\in C_{j}$, are set as positive keys $\{k_{+}\}$ and those from different clusters are set as negative $\{k_{-}\}=\mathbb{K}\backslash\{k_{+}\}$. In practice, we use bilinear products [38] for the score function in Eq. (3), and similarities between the query and keys are computed as $q^{\top}Wk$, where $W$ is a learnable parameter matrix. The InfoNCE loss in Eq. (3) is rearranged as
$\mathcal{L}_{K}\!=\!-\log\frac{\exp(q^{\top}Wk_{+})}{\exp(q^{\top}Wk_{+})+\exp(q^{\top}Wk_{-})}\!=\!-\log\frac{\sum\limits_{i^{\prime}\in C_{j}}\exp(z_{i}^{\top}Wz_{i^{\prime}})}{\sum\limits_{i^{\prime}\in C_{j}}\!\!\!\exp(z_{i}^{\top}Wz_{i^{\prime}})\!+\!\!\!\sum\limits_{i^{*}\notin C_{j}}\!\!\!\exp(z_{i}^{\top}Wz_{i^{*}})}.$ (4)
Following the MoCo method [42], we maintain a query encoder $\theta_{q}$ and a key encoder $\theta_{k}$, and use a momentum update to facilitate the key representations’ consistency as
@@ 2.2 Attention-Guided Role Coordination
After acquiring agents’ contrastive role representations from their local information, we introduce an attention mechanism in value decomposition to enhance agent coordination in the sophisticated role space with a global view. Popular CTDE algorithms, such as QMIX, realize behavior coordination across agents via a mixing network that estimates joint action-values as a monotonic combination of per-agent values, and the mixing network weights are conditioned on the system’s global state. Naturally, it is interesting to incorporate the learned role information into the mixing network to facilitate skillful coordination across roles. The simplest approach is to concatenate the global state and role representations for generating mixing network weights, while it fails to exploit the internal structure to effectively extract correlations in the role space. Fortunately, the attention mechanism [43] aligns perfectly with our intention by prompting the global state to attend to learned role patterns, thus providing more expressive credit assignment in value decomposition.
The attention mechanism aims to draw global dependencies without regard to their distance in the input or output sequences, and has gained substantial popularity as a fundamental building block of compelling sequence modeling and transduction models, such as GPT [44], vision transformers [45], and decision transformers [46]. An attention function can be described as mapping a query and a set of key-value pairs to a weighted sum of the values, where the weight assigned to each value is computed by a compatibility function of the query and corresponding key. As role representations are learned based on extracting agent behaviors from history trajectories, we also use a GRU to encode the history states $(\bm{s}^{0},\bm{s}^{1},...,\bm{s}^{t})$ into a state embedding $\bm{\tau}^{t}$ for facilitating information matching between states and role representations. Then, we set the state embedding $\bm{\tau}\in\mathbb{R}^{d_{s}\times d}$ as the query, and the role representations $\bm{z}=[z_{1},...,z_{n}]^{\top}\in\mathbb{R}^{n\times d}$ as both the key and value, where $d$ is the dimension of role representation and $d_{s}$ is the length of state embedding. Formally, we calculate a weighted combination of role representations as
@@ 3 Experiments
Implementations. We choose SMAC [35] as the first testbed for its rich maps and convenient visualization tools, and realize ACORM on top of the popular QMIX algorithm. For visualization, we render game scenes, and show agent embeddings and role representations using t-SNE. Appendix C gives evaluation results on the GRF benchmark, and Appendix D shows the algorithm architecture, experimental settings and results of MAPPO-based ACORM.
Baselines. We compare ACORM to QMIX and the five most relevant baselines: 1) RODE [29], a classical role-based method with action space decomposition; 2) EOI [20] that encourages diversified individuality via training an observation-to-identity classifier; 3) MACC [47] that uses attention to concentrate on most related subtasks; 4) CDS [19] that introduces diversity in both optimization and representation; 5) CIA [21] that boosts credit-level distinguishability via contrastive learning, and 6) GoMARL [34] that learns automatic grouping with dual-hierarchical value factorization.
@@ 4 Related Work
Agent Heterogeneity. As a compelling paradigm, the CTDE framework [7] has successfully yielded numerous algorithms [48, 49, 12]. Many of them share policy parameters to improve learning efficiency and scale to large-scale systems, which inevitably results in homogeneous behaviors across agents [31]. To promote diversity, SePS [18] partitions agents into a fixed set of groups and shares parameters within the same group only, while ignoring the evolving dynamics of the team. GoMARL [34] presents an automatic grouping mechanism and encourages diversity across different groups with a similarity-diversity regularizer. MAVEN [50] learns diverse exploratory behaviors by introducing a latent space for hierarchical control. CDS [19] equips each agent with an additional local Q-function to decompose the policy to the shared and non-shared part. EOI [20] promotes individuality by encouraging agents to visit their own familiar observations. CIA [21] boosts agent distinguishability in value decomposition via contrastive learning. While differentiating each agent from the rest, these methods neglect the development of effective team composition with implicit task allocation, and might hinder the discovery of sophisticated coordination patterns.
Role Emergence. Researchers have also introduced the role concept into multi-agent tasks [26, 27, 51, 52], or similarly, the concept of skills [53] or subtasks [47]. ROMA [28] conditions individual policies on roles and solely relies on the current observation to generate the role embedding, which might be inadequate for capturing complex agent behaviors. RODE [29] associates each role with a fixed subset of the full action space to reduce learning complexity. Following RODE, SIRD [54] transforms role discovery into hierarchical action space clustering. Nonetheless, they neglect the evolving dynamics of the team since roles are kept fixed in the training stage. For dynamic role assignment, some works learn identity representations to group agents during training, and maintain a selection strategy to realize the assignment from agents to skills [31] or subtasks [32, 33]. Nevertheless, they encode the identity solely from a one-hot vector, which might be insufficient to distinguish complex agent characteristics. COPA [30] realizes dynamic role allocation via periodically distributing a global view of team composition to each agent even in execution. However, it relaxes the CTDE constraint by introducing communication during decentralized execution, and the global composition is simply sampled from a fixed set of teams.
@@
Contrastive Learning and Attention Mechanism. Contrastive learning is gaining widespread popularity for self-supervised representation learning in various domains [42, 55, 38]. As a simple and effective technique, contrastive learning is also investigated to assist MARL tasks, such as boosting the credit-level distinguishability [21], facilitating the utilization of the agent-level contextual information [56], and grounding agent communication [57]. In this study, we apply contrastive learning to optimize role representations, facilitating sophisticated coordination with better role assignment.
Attention is the fundamental building block of famous transformer architectures [43] that exhibit growing dominance in advances of AI research [44, 45, 46]. Due to its superiority in extracting dependencies between sequences, attention has been widely applied to MARL domains for various utilities, such as learning a centralized critic [15], concentrating on relevant subtasks [47], formulating MARL as a sequence modeling problem [58], addressing stochastic partial observability [59], etc [60, 61, 62]. In this study, we use attention to guide skillful role coordination for more expressive credit assignment.

### 02 Mar 2024 v2

@@ Abstract
Real-world multi-agent tasks usually involve dynamic team composition with the emergence of roles, which should also be a key to efficient cooperation in multi-agent reinforcement learning (MARL). Drawing inspiration from the correlation between roles and agent’s behavior patterns, we propose a novel framework of Attention-guided COntrastive Role representation learning for MARL (ACORM) to promote behavior heterogeneity, knowledge transfer, and skillful coordination across agents. First, we introduce mutual information maximization to formalize role representation learning, derive a contrastive learning objective, and concisely approximate the distribution of negative pairs. Second, we leverage an attention mechanism to prompt the global state to attend to learned role representations in value decomposition, implicitly guiding agent coordination in a skillful role space to yield more expressive credit assignment. Experiments on challenging StarCraft II micromanagement and Google research football tasks demonstrate the state-of-the-art performance of our method and its advantages over existing approaches. Our code is available at https://github.com/NJU-RL/ACORM.
@@ 1 INTRODUCTION
Cooperative multi-agent reinforcement learning (MARL) aims to coordinate a system of agents towards optimizing global returns (Vinyals et al., 2019), and has witnessed significant prospects in various domains, such as autonomous vehicles (Zhou et al., 2020), smart grid (Chen et al., 2021a), robotics (Yu et al., 2023), and social science (Leibo et al., 2017). Training reliable control policies for coordinating such systems remains a major challenge. The centralized training with decentralized execution (CTDE) (Foerster et al., 2016) hybrids the merits of independent Q-learning (Foerster et al., 2017) and joint action learning (Sukhbaatar et al., 2016), and becomes a compelling paradigm that exploits the centralized training opportunity for training fully decentralized policies (Wang et al., 2023). Subsequently, numerous popular algorithms are proposed, including VDN (Sunehag et al., 2018), QMIX (Rashid et al., 2020), MAAC (Iqbal & Sha, 2019), and MAPPO (Yu et al., 2022).
Sharing policy parameters is crucial for scaling these algorithms to massive agents with accelerated cooperation learning (Fu et al., 2022). However, it is widely observed that agents tend to acquire homogeneous behaviors, which might hinder diversified exploration and sophisticated coordination (Christianos et al., 2021). Some methods (Li et al., 2021; Jiang & Lu, 2021; Liu et al., 2023) attempt to promote individualized behaviors by distinguishing each agent from the others, while they often neglect the prospect of effective team composition with implicit task allocation. Real-world multi-agent tasks usually involve dynamic team composition with the emergence of roles (Shao et al., 2022; Hu et al., 2022) [^1] Taking the football game (Kurach et al., 2020) as an example, the midfielders are primarily responsible for delivering the ball to the forwards to coordinate shots on goal in the offensive phase, while they need to drop back and join the defenders to block passing lanes on the defensive. Early works introduce the role concept into multi-agent systems (Dastani et al., 2003; Sims et al., 2008; Lhaksmana et al., 2018), while they usually require prior domain knowledge to pre-define role responsibilities. Recently, ROMA (Wang et al., 2020) learns emergent roles conditioned solely on current observations, and RODE (Wang et al., 2021) associates each role with a fixed subset of the joint action space. COPA (Liu et al., 2021) allows dynamic role allocation via distributing a global view of team composition to each agent during execution. Some works decompose the task into a set of skills (Liu et al., 2022) or subtasks (Yang et al., 2022; Iqbal et al., 2022) with a hierarchical structure for control. Overall, existing role-based methods still suffer from several deficiencies, such as insufficient characterization of complex behaviors for role emergence, neglect of evolving team dynamics, or relaxation of the CTDE constraint.
@@
ACORM is fully compatible with CTDE methods, and we realize ACORM on top of two popular MARL algorithms, QMIX (Rashid et al., 2020) and MAPPO (Yu et al., 2022), benchmarked on challenging StarCraft multi-agent challenge (SMAC) (Samvelyan et al., 2019) and Google research football (GRF) (Kurach et al., 2020) environments. Experiments demonstrate that ACORM achieves state-of-the-art performance on most scenarios. Visualizations of learned role representations, heterogeneous behavior patterns, and attentional value decomposition shed further light on our advantages. Ablation studies confirm that ACORM promotes higher coordination capacity by virtue of contrastive role representation learning and attention-guided credit assignment, respectively, even if agents have the same innate characteristics. In summary, our contributions are threefold:
@@ 2 Method
In this section, we present the ACORM framework. We consider cooperative multi-agent tasks formulated as a Dec-POMDP (Oliehoek & Amato, 2016), $G=\langle I,S,A,P,R,\Omega,O,n,\gamma\rangle$, where $I$ is a finite set of $n$ agents, $s\in S$ is the global state, and $\gamma\in[0,1)$ is the discount factor. At each time step, each agent $i$ draws an observation $o_{i}\in O$ from $\Omega(s,i)$ and selects a local action $a_{i}\in A$. After executing the joint action $\bm{a}=[a_{1},...,a_{n}]^{\top}\in A^{n}$, the system transitions to a next state $s^{\prime}$ according to $P(s^{\prime}|s,\bm{a})$ and receives a reward $r=R(s,\bm{a})$ shared by all agents.
@@
ACORM consists of individual Q-networks in Fig. 1(b) and a mixing network in Fig. 1(a). We introduce mutual information maximization to formalize role representation learning, and derive a contrastive learning objective that optimizes agent embeddings $\{e^{t}_{i}\}_{i=1}^{n}$ in a self-supervised way to acquire contrastive role representations $\{z_{i}^{t}\}_{i=1}^{n}$, which is shown in Fig. 1(c) and will be introduced in detail in Section 2.1. In value decomposition, we employ multi-head attention (MHA) to prompt the global state to attend to learned role patterns, guiding skillful agent coordination in the high-level role space for facilitating expressive credit assignment, as described in Fig. 1(d) and in Section 2.2. Appendix B presents the pseudocode, and Appendix C gives the extension to MAPPO.
@@ 2.1 Contrastive Role Representations
In practice, directly optimizing mutual information is intractable. Inspired by the noise contrastive estimation (InfoNCE) (Oord et al., 2018) in the literature of contrastive learning (Laskin et al., 2020; Yuan & Lu, 2022), we derive a lower bound of Eq. (1) with the following theorem.
@@
Negative Pairs Generation. Periodically, we partition all $n$ agents into $K$ clusters $\{C_{j}\}_{j=1}^{K}$ according to agent embeddings. [^4] In this paper, we simply use K-means (Hartigan & Wong, 1979) based on Euclidean distances between agent embeddings. Moreover, it can be easily extended to more complex clustering methods such as Gaussian mixture models (Bishop, 2006). In Appendix D, we conduct the hyperparameter analysis about the influence of different $K$ values and how to determine the number of clusters automatically. Naturally, we encourage role representations from the same cluster to stay close to each other, while differing from agents in other clusters. For agent $i$, we denote its role representation $z_{i}$ as the query $q$, and role representations of other agents as the keys $\mathbb{K}=\{z_{1},...,z_{n}\}\backslash z_{i}$. Points from the same cluster as the query, $i\in C_{j}$, are set as positive keys $\{k_{+}\}$ and those from different clusters are set as negative $\{k_{-}\}=\mathbb{K}\backslash\{k_{+}\}$. In practice, we use bilinear products (Laskin et al., 2020) for the score function in Eq. (3), and similarities between the query and keys are computed as $q^{\top}Wk$, where $W$ is a learnable parameter matrix. The InfoNCE loss in Eq. (3) is rearranged as
$\mathcal{L}_{K}\!=\!-\log\frac{\exp(q^{\top}Wk_{+})}{\exp(q^{\top}Wk_{+})+\exp(q^{\top}Wk_{-})}\!=\!-\log\frac{\sum\nolimits_{i^{\prime}\in C_{j}}\exp(z_{i}^{\top}Wz_{i^{\prime}})}{\sum\limits_{i^{\prime}\in C_{j}}\!\!\!\exp(z_{i}^{\top}Wz_{i^{\prime}})\!+\!\!\!\sum\limits_{i^{*}\notin C_{j}}\!\!\!\exp(z_{i}^{\top}Wz_{i^{*}})}.$ (4)
Following the MoCo method (He et al., 2020), we maintain a query encoder $\theta_{q}$ and a key encoder $\theta_{k}$, and use a momentum update to facilitate the key representations’ consistency as
@@ 2.2 Attention-Guided Role Coordination
After acquiring agents’ contrastive role representations from their local information, we introduce an attention mechanism in value decomposition to enhance agent coordination in the sophisticated role space with a global view. Popular CTDE algorithms, such as QMIX, realize behavior coordination across agents via a mixing network that estimates joint action-values as a monotonic combination of per-agent values, and the mixing network weights are conditioned on the system’s global state. Naturally, it is interesting to incorporate the learned role information into the mixing network to facilitate skillful coordination across roles. The simplest approach is to concatenate the global state and role representations for generating mixing network weights, while it fails to exploit the internal structure to effectively extract correlations in the role space. Fortunately, the attention mechanism (Vaswani et al., 2017) aligns perfectly with our intention by prompting the global state to attend to learned role patterns, thus providing more expressive credit assignment in value decomposition.
The attention mechanism aims to draw global dependencies without regard to their distance in the input or output sequences, and has gained substantial popularity as a fundamental building block of compelling sequence modeling and transduction models, such as GPT (Brown et al., 2020), vision transformers (Dosovitskiy et al., 2021), and decision transformers (Chen et al., 2021b). An attention function can be described as mapping a query and a set of key-value pairs to a weighted sum of the values, where the weight assigned to each value is computed by a compatibility function of the query and corresponding key. As role representations are learned based on extracting agent behaviors from history trajectories, we also use a GRU to encode the history states $(\bm{s}^{0},\bm{s}^{1},...,\bm{s}^{t})$ into a state embedding $\bm{\tau}^{t}$ for facilitating information matching between states and role representations. Then, we set the state embedding $\bm{\tau}\in\mathbb{R}^{d_{s}\times d}$ as the query, and the role representations $\bm{z}=[z_{1},...,z_{n}]^{\top}\in\mathbb{R}^{n\times d}$ as both the key and value, where $d$ is the dimension of role representation and $d_{s}$ is the length of state embedding. Formally, we calculate a weighted combination of role representations as
@@ 3 Experiments
Implementations. We choose SMAC (Samvelyan et al., 2019) as the first testbed for its rich maps and convenient visualization tools, and realize ACORM on top of the popular QMIX algorithm. For visualization, we render game scenes, and show agent embeddings and role representations using t-SNE. Appendix C gives evaluation results on the GRF benchmark, and Appendix D shows the algorithm architecture, experimental settings and results of MAPPO-based ACORM.
Baselines. We compare ACORM to QMIX and six baselines: 1) RODE (Wang et al., 2021) with action space decomposition; 2) EOI (Jiang & Lu, 2021) that encourages diversified individuality via training an observation-to-identity classifier; 3) MACC (Yuan et al., 2022) that uses attention to concentrate on most related subtasks; 4) CDS (Li et al., 2021) that introduces diversity in both optimization and representation; 5) CIA (Liu et al., 2023) that boosts credit-level distinguishability via contrastive learning; and 6) GoMARL (Zang et al., 2023) with an automatic grouping mechanism.
@@ 4 Related Work
Agent Heterogeneity. As a compelling paradigm, CTDE (Foerster et al., 2016) has yielded numerous algorithms (Lowe et al., 2017; Son et al., 2019; Wang et al., 2023). Many of them share policy parameters to improve learning efficiency and scale to large-scale systems, which results in homogeneous behaviors across agents (Liu et al., 2022). To promote diversity, SePS (Christianos et al., 2021) partitions agents into a fixed set of groups and shares parameters within the same group only, while ignoring evolving dynamics of the team. GoMARL (Zang et al., 2023) generates dynamic groups with an automatic grouping mechanism to possess diverse strategies. MAVEN (Mahajan et al., 2019) learns diverse exploratory behaviors by introducing a latent space for hierarchical control. CDS (Li et al., 2021) equips each agent with an additional local Q-function to decompose the policy to the shared and non-shared part. EOI (Jiang & Lu, 2021) promotes individuality by encouraging agents to visit their own familiar observations. CIA (Liu et al., 2023) boosts agent distinguishability in value decomposition via contrastive learning. While differentiating each agent from the rest, these methods neglect the development of effective team composition with implicit task allocation, and might hinder the discovery of sophisticated coordination patterns.
Role Emergence. Researchers have also introduced the role concept into multi-agent tasks (Sims et al., 2008; Lhaksmana et al., 2018; Xia et al., 2023; Cao et al., 2023), or similarly, the concept of skills (Yang et al., 2020a) or subtasks (Yuan et al., 2022). ROMA (Wang et al., 2020) conditions individual policies on roles and solely relies on the current observation to generate the role embedding, which might be inadequate for capturing complex agent behaviors. RODE (Wang et al., 2021) associates each role with a fixed subset of the full action space to reduce learning complexity. Following RODE, SIRD (Zeng et al., 2023) transforms role discovery into hierarchical action space clustering. Nonetheless, they neglect the evolving dynamics of the team since roles are kept fixed in the training stage. For dynamic role assignment, some works learn identity representations to group agents during training, and maintain a selection strategy to realize the assignment from agents to skills (Liu et al., 2022) or subtasks (Yang et al., 2022; Iqbal et al., 2022). Nevertheless, they encode the identity solely from a one-hot vector, which might be insufficient to distinguish complex agent characteristics. COPA (Liu et al., 2021) realizes dynamic role allocation via periodically distributing a global view of team composition to each agent even in execution. However, it relaxes the CTDE constraint by introducing communication during decentralized execution, and the global composition is simply sampled from a fixed set of teams.
@@
Contrastive Learning and Attention Mechanism. Contrastive learning is gaining widespread popularity for self-supervised representation learning in various domains (He et al., 2020; Su et al., 2022; Laskin et al., 2020). As a simple and effective technique, contrastive learning is also investigated to assist MARL tasks, such as boosting the credit-level distinguishability (Liu et al., 2023), facilitating the utilization of the agent-level contextual information (Song et al., 2023), and grounding agent communication (Lo & Sengupta, 2022). In this study, we apply contrastive learning to optimize role representations, facilitating sophisticated coordination with better role assignment.
Attention is the fundamental building block of famous transformer architectures (Vaswani et al., 2017) that exhibit growing dominance in advances of AI research (Brown et al., 2020; Dosovitskiy et al., 2021; Chen et al., 2021b). Due to its superiority in extracting dependencies between sequences, attention has been widely applied to MARL domains for various utilities, such as learning a centralized critic (Iqbal & Sha, 2019), concentrating on relevant subtasks (Yuan et al., 2022), formulating MARL as a sequence modeling problem (Wen et al., 2022), addressing stochastic partial observability (Phan et al., 2023), etc (Yang et al., 2020b; Shao et al., 2023; Zhai et al., 2023). In this study, we use attention to guide skillful role coordination for more expressive credit assignment.